# Achieving Pareto Optimality in Games via Single-bit Feedback

### v1

@@ 5 Simulation
Figure 1: Performance of Algorithm 1 with $n=10$ agents and $m=2$ actions. The x-axis shows the exploration length $K$ (log scale), and the y-axis the total utility. Solid lines denote the exploited profile $a^{*}$, dashed lines the most frequently observed joint actions, and the black dashed line the socially optimal utility.
Figure 2: Limit values of $\epsilon$ as a function of the welfare gap $\Delta_{1}$ for $n=10$ agents and $m=2$ actions. The blue curve indicates the mean of the limiting $\epsilon$ values, while the orange curve shows the minimum acceptable $\epsilon$.
We evaluate the proposed Single-Bit Coordination Dynamics for Pareto-Efficient Outcomes (SBC-PE) through simulations. The system has $n=10$ agents, each with $m=2$ actions and with parameters $w_{i}=1,\lambda_{i}=0.2\ \forall i\in N$. Utilities $u_{i}(a)$ are generated randomly between 0 and 1, so they may not satisfy the minimum utility constraints (1).
Fig. 1 shows the effect of the exploration horizon $K$ on aggregate welfare. Larger $K$ improves reliability of convergence to the Pareto-efficient profile within $10^{7}$ iterations, while large $\epsilon$ values (e.g., $\epsilon=0.5$) hinder stability.
Fig. 2 shows limiting $\epsilon$ values as a function of the welfare gap $\Delta_{1}$, defined as the difference between the best and second-best aggregate utilities. As $\Delta_{1}$ widens, the frequency of the socially maximizing action increases, allowing larger $\epsilon$ values and highlighting the robustness of SBC-PE when the optimal solution is well separated from suboptimal ones.
@@ 6 Discussion & Conclusion
Simulations confirm that SBC-PE achieves exact socially optimal outcomes with minimal communication. The parameter $\epsilon$ is critical: too small slows adaptation, while too large reduces stability. The exploration horizon $K$ must also scale with system size to ensure reliable identification of Pareto-efficient profiles.
Remark 4. If agents can observe the full joint action, the message rule in (3) modifies to
$m_{i}(t)=\begin{cases}\mathbb{I}_{\{u_{i}(t)>\lambda_{i},\,a_{i}(t)\in BR_{i}(a_{-i}(t))\}}&\text{w.p. }\epsilon^{\,1-w_{i}u_{i}(t)},\\ 0&\text{otherwise},\end{cases}$
where $BR_{i}(a_{-i}(t))$ is the best-response set of agent $i$. Under this rule, if a pure equilibrium exists, Algorithm 1 converges to a Pareto Optimal Equilibrium, even for relatively large $\epsilon$.
Communication-based methods guaranteeing Pareto optimality often face scalability challenges from heavy message exchange. SBC-PE remains efficient since each agent broadcasts only one bit per iteration, independent of $n$ or $m$. Thus, the overall communication cost grows linearly with the number of agents while message size is fixed. Hence, the algorithm is communication efficient and scalable. By combining exact convergence guarantees with minimal signaling requirements, SBC-PE demonstrates that strong welfare optimization is possible even under very limited communication. This makes it a practical candidate for large-scale applications.

### v2

@@ 5 Simulation
Figure 1: Performance of Algorithm 1 for $n=10$ agents. The x-axis is the exploration length $K$ (log scale), and the y-axis the total utility. Solid lines depict the committed joint action $\bar{a}$, dashed lines the most-frequently-played content-endorsed joint action, and the black dotted line the optimal utility.
Figure 2: Limit values of $\epsilon$ as a function of the welfare gap $\Delta_{1}$ for $n=10$ agents and two actions. The blue curve indicates the mean of the limiting $\epsilon$ values, while the orange curve shows the minimum acceptable $\epsilon$.
We examine the performance of the proposed SBC-PE algorithm for different exploration lengths $K$, $\epsilon$ parameters and the welfare gap $\Delta_{1}$, as described in (7). To this end, we focus on $(n=10)$-agent games in which each agent $i\in N$ has two actions, $w_{i}=1$, and $\lambda_{i}=0.2$. We generate utilities $u_{i}(a)$ randomly between 0 and $1$ so they may not satisfy the minimum utility constraints (2).
In Fig. 1, we examine the impact of $K$ and $\epsilon$ while comparing the performances of committed joint action and the most-frequently-played content-endorsed action, i.e., $\operatorname*{arg\,max}_{a}\widehat{\theta}(a)$. This illustrates that larger $\epsilon$ values can hinder efficiency while smaller $\epsilon$ values can require longer exploration. Note that Assumption 1-(iii) provides an explicit bound on $\epsilon$ ensuring Pareto-efficient outcomes for sufficiently large $K$, as shown in Theorem 4.
In Fig. 2, we further examine the average and the minimum $\epsilon$ parameters ensuring Pareto-efficient outcomes across 20,00020{,}000 game samples for different welfare gaps $\Delta_{1}$. For large $\Delta_{1}$, the frequency of the content-endorsed socially maximizing action increases (see Remark 4). This allows larger $\epsilon$ values, showing the robustness of SBC-PE when the optimal solution is well separated from suboptimal ones.
@@ 6 Discussion & Conclusion
Simulations confirm that SBC-PE achieves exact socially optimal outcomes with minimal communication. The parameter $\epsilon$ is critical: too small leads to slow adaptation, while too large reduces efficiency. The exploration horizon $K$ must also scale with system size to ensure reliable identification of Pareto-efficient profiles.
Remark 4. Consider that agents can observe the full joint action and use the message rule
$m_{i}^{t}=\begin{cases}\mathbb{I}_{\{a_{i}^{t}\in\mathrm{BR}_{i}(a_{-i}^{t})\}}&\textrm{with probability }\epsilon^{\,1-w_{i}\cdot u_{i}^{t}},\\ 0&\textrm{otherwise},\end{cases}$
where $\mathrm{BR}_{i}(a_{-i}^{t})\coloneqq\operatorname*{arg\,max}_{a_{i}\in A^{i}}\{u^{i}(a_{i},a_{-i}^{t})\}$ is the best response set. Then, if a pure equilibrium exists, Algorithm 1 can learn Pareto-Optimal Equilibrium, even for relatively large $\epsilon$.
Communication-based methods guaranteeing Pareto optimality often face scalability challenges from heavy message exchange. SBC-PE remains efficient since each agent broadcasts only one bit per stage, independent of the game size. Thus, the overall communication cost grows linearly with the number of agents while message size is fixed. Hence, the algorithm is communication efficient and scalable. SBC-PE demonstrates that strong welfare optimization is possible even under very limited communication. This makes it a practical candidate for large-scale applications.